# Efficient Event-based Semantic Segmentation with Spike-driven Lightweight Transformer-based Networks

### v1

@@ IV-E3 Loss Design
V CONCLUSIONS
In this work, we tackle the real-world challenges of traditional semantic segmentation by designing lightweight single-branch architecture and integrating them into the spike-based paradigm. We design a spike-driven lightweight transformer-based network, named SLTNet, that is built on efficient SCBs and STBs basic blocks. Extensive experimental results show that our work outperforms SOTA SNN-based methods on both DDD17 and DSEC-Semantic datasets. SLTNet also achieves 2.01$\times$ more computational load saving, 5.48$\times$ more energy-saving, and 1.14$\times$ faster inference speed. In the future, we will extend our method to other fields, such as simultaneous localization mapping, and flow estimation.

### v2

@@ V-E3 Loss Design
VI Conclusion
In this work, we tackle the real-world challenges of traditional semantic segmentation by designing a lightweight single-branch architecture and integrating it into the spike-based event segmentation paradigm. Specifically, we design a SLTNet that is built on novel SCBs and STBs basic blocks, where SCBs are designed for efficient feature extraction, while STBs are proposed to achieve long-range contextual feature interaction. Extensive experimental results show that our work outperforms SOTA SNN-based methods on both DDD17 and DSEC-Semantic datasets. SLTNet also achieves 4.58$\times$ more energy-saving, and 114 FPS faster inference speed.
Given the accuracy gap between ANN-based methods and SLTNet, future work will focus on developing a more advanced SNN architecture for event processing and extending our approach to other domains, such as autonomous driving and robotic localization.